OpenStudy (johnnydicamillo):

find f'(x) when f(x) = ln(1/x)

OpenStudy (anonymous):

\[\frac{d}{dx}\left[\ln\frac{1}{x}\right]=\frac{\dfrac{d}{dx}\left[\dfrac{1}{x}\right]}{\dfrac{1}{x}}\]